Information pertaining to Yekstop Corp.'s sales revenue is presented below: November December January Cash sales $ 100,000 $ 129,000 $ 82,000 Credit sales 292,000 454,000 238,000 Total sales $ 392,000 $ 583,000 $ 320,000 Management estimates that 5% of credit sales are eventually uncollectible. Of the collectible credit sales, 65% are likely to be collected in the month of sale and the remainder in the month following the month of sale. The company desires to begin each month with an inventory equal to 75% of the sales projected for the month. All purchases of inventory are on open account; 20% will be paid in the month of purchase, and the remainder paid in the month following the month of purchase. Purchase costs are approximately 60% of the selling prices. Budgeted January cash payments for December inventory purchases by Yekstop Corp. are:

Multiple Choice

$46,290.

$69,960.

$185,160.

$279,840.

$292,075.
Answer #1

December purchase = 583000*60% = 349800

January Cash payment for December inventory purchase = 349800*80% = 279840

So answer is d) $279840